OpenStudy (anonymous):

To solve by completing the square, what value should you add to each side of the equation? x^2+10x= -4

OpenStudy (raden):

add by 1/2 of the coefficient of x, then squares so, u have to add by (10/2)^2 = 5^2 = 25 to both sides